OpenStudy (anonymous):

sec x sin (csc x + cos x) / sec x + sin x ... what is the answer simplified (number)

OpenStudy (anonymous):

\[\frac{ (\frac{ 1 }{ cosx }. \sin x)(\frac{ 1 }{\sin x }+ cosx) }{ \frac{ 1 }{ \cos x} + \sin x } = 1\]